# Threshold Graphs are Globally Synchronizing

### v1

@@ Abstract.
Place $n$ nodes on the unit circle and connect some pairs with springs that pull connected nodes toward each other. Which graphs always drive all nodes to converge to a single point, no matter where they start? This is known as the global synchronization problem of the Kuramoto model and is closely tied to the benign nonconvexity of its synchronization landscape. In this work, we show how this dynamical and continuous question can be reduced to a static and combinatorial one. We prove that threshold graphs, the graph class closed under adding isolated nodes and adding universal nodes, are globally synchronizing, based purely on a combinatorial analysis of unit-vector configurations in the plane. It is particularly striking that threshold graphs range from extremely sparse structures such as stars to the complete graph, suggesting that global synchronization in this class cannot be attributed to density. Our proof shows that global synchronization can instead emerge from a form of geometric symmetry.
@@ 1. Introduction
The study of synchronization traces back to the 17th century, when Christiaan Huygens first observed the spontaneous synchronization of pendulum clocks. Since then, synchronization phenomena have been widely studied across both natural and technological systems, from the flashing of fireflies [BB68, SS93] and the rhythmic beating of heart cells [Win67, MS06] to synchronization in power grids [DB13], and the Transformer architecture [GLPR23a, CRMB24]. The Kuramoto model, introduced by Yoshiki Kuramoto in 1975 [Kur75], strikes a balance between mathematical tractability and sufficient complexity, has become one of the key mathematical models for understanding the collective dynamics of coupled oscillators.
The Kuramoto model describes a system of $n$ oscillators continuously evolving on the unit circle. Each vertex $i\in V$ corresponds to an oscillator with phase (angle) $\theta_{i}:\mathbb{R}\rightarrow\mathbb{S}^{1}$. The dynamics of phase $\theta_{i}$ of each oscillator evolves according to:
$\frac{d\theta_{i}}{dt}=\omega_{i}+\sum_{j=1}^{n}\boldsymbol{A}_{ij}\sin(\theta_{j}-\theta_{i}),$ (1)
This means that each oscillator’s phase dynamics is governed by its intrinsic frequency $\omega_{i}$ while also being influenced by the phase differences with its neighbors in a nonlinear manner. There are two lines of research regarding this model focusing heterogenous and homogenouse case respectively.
Studies on the former case primarily consider a systems of oscillators with heterogenous intrinsic frequences $\{\omega_{i}\}_{i=1}^{N}$ coupled through a complete graph with identical weight $\boldsymbol{A}_{ij}=\frac{K}{N}$. Here, $K$ determines how strongly oscillators influence each other. A fundamental question in the heterogeneous setting is to determine the critical coupling strength $K_{c}$ such that, for $K>K_{c}$, the oscillators achieve frequency synchronization, i.e., $\dot{\theta}_{i}-\dot{\theta}_{j}\to 0$, meaning that all oscillators rotate with the same angular velocity. If this indeed holds, it is natural to further ask whether the oscillators not only rotate with the same angular velocity, but also converge to the same phase, known as phase synchronization, that is, whether $\theta_{i}-\theta_{j}\rightarrow 0$, as $t\rightarrow\infty$. The study of synchronization in the heterogeneous Kuramoto model is particularly relevant to real-world systems. In many natural and engineered networks such as power grids, neuronal circuits, and circadian rhythms, the individual oscillators typically have distinct natural frequencies. Within this heterogeneous setting, a rich variety of nonlinear dynamical phenomena can emerge, including partial synchronization, multistability, and even chaotic behavior. For further background and details, we refer the reader to [Str00, ABV+05, DB14].
Recent works on the Kuramoto model has moved beyond the fully connected case to more general interaction graphs. Such settings are not only mathematically appealing but also more realistic, since interactions in natural and engineered systems are rarely all-to-all. In this work, we focus on the homogeneous case with identical intrinsic frequencies ($\omega_{i}=\omega$ for all $i$) and aim to understand how graph structure governs synchronization. To begin with, we first remark that in this case, frequency synchronization always holds and the remaining question is whether phasr synchronization can be achieved. This is not obvious if we only look at the equation $\dot{\theta}_{i}=\omega+\sum_{j=1}^{n}\boldsymbol{A}_{ij}\sin(\theta_{j}-\theta_{i})$, for all $i$, since it tells that the instantaneous angular velocity of each node depends both on this common intrinsic term and on the coupling with its neighbors. However, applying a co-rotating frame $\theta_{i}(t)\leftarrow\theta_{i}(t)-\omega t$, the equations becomes
$\frac{d\theta_{i}}{dt}=\sum_{j=1}^{n}\boldsymbol{A}_{ij}\sin(\theta_{j}-\theta_{i}),\quad i\in\{1,\cdots,n\},$ (2)
and this reduced system is in fact a gradient flow of the Kuramoto energy function
$E(\theta)=\frac{1}{2}\sum_{1\leq i<j\leq n}\boldsymbol{A}_{ij}(1-\cos(\theta_{i}-\theta_{j}))$ (3)
as its gradient satisfies $\dot{\theta_{i}}{dt}=-\frac{\partial E}{\partial\theta_{i}}$. By LaSalle’s invariance principle, every trajectory converges to the invariant set of equilibria satisfying $\dot{\theta}_{i}=0$ for all $i$. Hence, the angle differences do not lead to frequency discrepancies in the asymptotic sense. In other words, frequency synchronization always holds, as they share common frequency $\omega$. Now the real challenge in the homogeneous case lies in understanding whether phase synchronization can always be achieved, regardless of the initial phase configurations. It is evident that this property is entirely determined by the graph structure, since the adjacency matrix $\mathbf{A}$ of the graph is the only external parameter in the homogeneous Kuramoto model (2) and equivalently in its energy formulation (3). Thus, we define:
@@
We say that a graph $G$ with adjacency matrix $\mathbf{A}$ is globally synchronizing if, all oscillators converge to the same phase, i.e.,
$\lim_{t\to\infty}\theta_{i}(t)-\theta_{j}(t)=0\quad\text{for all }i,j\in V.$
for all initial conditions $\boldsymbol{\theta}(0)\in\mathbb{R}^{n}$ except for a set of measure zero. In other words, almost all trajectory converges to the synchronous state $\theta_{i}=\theta_{j}$, for all $i,j\in V(G)$.
@@
On which graphs, the Kuramoto model (2) exhibits global synchronization?
@@
This seemingly simple description conceals a mathematically challenging task. To address this problem, we note that the Kuramoto model (2) can be formulated as a gradient flow of the energy function (3), as mentioned before. Hence, understanding the geometry of (3) provides direct insight into the long-term dynamics of (2). In particular, the synchronous states correspond exactly to the global minima of this energy function, and those minima that are local but not global (called spurious local minimua) correspond to non-synchronous equilibria. If such spurious local minima exist, the gradient dynamics may converge to them whenever the initialization lies within their basin of attraction, in which case the graph $G$ is not globally synchronizing. In contrast, if the energy function admits no spurious local minima, then the graph $G$ is globally synchronizing, since this rules out convergence to any non-synchronous equilibrium except possibly for a set of initial conditions of measure zero. A rigorous proof of this statement can be established using the Łojasiewicz gradient inequality together with the center manifold theorem; see Lemma A.1 in [GLPR23b] for an excellent exposition.
Contributions and prior work. Aiming to understand which graphs and how them foster global synchronization, previous studies on this problem can be broadly divided into two lines of research: (i) investigating whether random graphs globally synchronize, and (ii) determining the minimum edge density required to ensure global synchronization. Both approaches are, in a certain sense, grounded in the fact that the complete graph is globally synchronizing, and that both random graphs and very dense graphs tend to behave like complete graphs. Intuitively, a graph of sufficiently high density becomes homogeneous enough to approximate a complete graph. In this sense, full connectivity is not strictly required, that is, global synchronization can still emerge with a small gap from the fully connected case.
In this work, we prove that threshold graphs, as a graph class that by no means closed to complete, but still globally synchronize.
(i) What are threshold graphs? Threshold graphs can be constructed starting from a single vertex by iteratively adding each new vertex as either: an isolated vertex connecting to no existing vertex, or a dominating vertex connecting to all existing vertices. This generative process is uniquely encoded by a binary sequence of length n1n-1, where 0 denotes an isolated vertex and 1 a dominating one. Thus our goal is to show all threshold graphs with sequence ending by $1$ must globally synchronizes, because otherwise it is not connected.
Figure 1. Four connected threshold graphs on 1919 vertices, from the sparsest ($S_{18}$) to the densest ($K_{19}$). Node labels show the construction order: 11 is the initial vertex, later numbers follow the binary code. White and black circles denote isolated and dominating vertices (bit 0 and bit 11). Codes from left to right: 000000000000000001, 000100000000010001, 010001101101010001, $111111111$.
(ii) Why threshold graphs? Many theorems start from initial observations, from which suitable conjectures are formulated. Our case is no exception. Our investigation into global synchronization on threshold graphs began with a simple observation: complete graphs and star graphs, respectively the densest and sparsest connected graphs in the class of threshold graphs, both exhibit global synchronization. Certainly, this is not sufficient to propose the conjecture that all threshold graphs globally synchronize. Indeed, although the two extremes synchronize, this does not necessarily imply that the entire spectrum of threshold graphs does.
However, two further observations strengthened our confidence in this conjecture. Firstly, in threshold graphs, the neighborhoods of any two nodes are nested: for any pair $u,v$, either $N[u]\subseteq N(v)$ or $N[v]\subseteq N(u)$. This total preorder of vicinities suggests that influence may propagate step by step from the most dominant node to the rest without interference, hinting at the possibility of global synchronization. Secondly, no counter-example was found in our numerical experiments up to 50 size graphs [^1] First, we transform the equilibrium equations of the Kuramoto model into a system of polynomial equations. We then use Bertini software package [BHSW13] to compute the zeros of this system, and subsequently evaluate the Jacobian at each solution to determine their stability.. These evidence encourages us to investigate whether all connected threshold graphs global synchronize. In particular, we established the following result:
@@
Many existing proofs of global synchronization rely, either explicitly or implicitly, on the so-called half-circle lemma. This lemma states that if all oscillators are confined to an open half-circle, then their dynamics will contract toward synchrony. Based on this observation, the bootstrap strategy attempts to show that all second-order critical points must lie within a half-circle. If successful, this rules out the existence of non-synchronous equilibria and thereby implies global synchronization.
@@
In this work, we bypass the half-circle approach by reducing the high-dimensional geometric question of whether the energy landscape is benign to a planar vector-geometry condition. This condition naturally leads us to study structural twins (pairs of nodes sharing the same neighborhood in the graph) as well as a new notion we introduce, that of geometric twins (a broader concept depending jointly on the graph structure and the equilibrium configuration $\theta$). These notions clarify how the relative positions of phasors behave at second order stationary point of the energy landscape $\mathbb{E}(\boldsymbol{\theta)}$. Building on this understanding, we employ an inductive strategy: synchronization is first established locally on small substructures, and then shown to propagate across the whole graph along the neighborhood-inclusion hierarchy of threshold graphs.
(iii) What we learned from global synchrony of threshold graphs? While identifying previously unknown classes of globally synchronizing graphs contributes to the expansion of known results, a more fundamental objective of this line of research is to elucidate the mechanisms that give rise to global synchronization itself. It is already known that for certain classes of graphs, global synchronization arises through mechanisms that, in a certain sense, are close to those of complete graphs. Threshold graphs, however, lie outside this regime. Nevertheless, we prove in this work that they also exhibit global synchronization. What deeper insight can we draw from this result?
Our analysis shows that the synchronization of threshold graphs is instead driven by another mechanism, namely their intrinsic symmetries. These symmetries are not merely combinatorial, such as sets of nodes sharing identical neighborhoods, but are geometric in nature. We will introduce this notion in detail later, but to give an intuition, these geometric symmetries emerge inductively as the graph is reduced node by node. Through this process, we uncover certain sets of nodes that, while not symmetric in the structural sense of graph theory, exhibit geometric symmetry at the second-order critical points of (3), which allows us to show that these nodes form a synchronized group. As the reduction continues, more geometric symmetries are revealed, leading to the formation of progressively larger synchronized groups, until the entire node set forms a single synchronized group, thereby guaranteeing that all second order critical points are synchronized states, thus the graph is global synchronizating.
1.1. Notation
@@
Throughout this work, we consider only connected simple graphs, i.e., graphs without self-loops or weighted edges. Furthermore, since the homogeneous Kuramoto model (2) on $G$ is fully determined by the graph topology, we sometimes use system $G$ to denote the Kuramoto model on this graph for simplicity. Following the convention of dynamical systems, we refer to the variable $\theta=(\theta_{1},\cdots,\theta_{n})$ of the Kuramoto model (2) as a state, and call the solution $\boldsymbol{\theta}(t)$ with respect to the initial condition $\boldsymbol{\theta}(0)$ in the Kuramoto model (2) a trajectory starting from $\boldsymbol{\theta}(0)$.

### v2

@@ Abstract.
The Kuramoto model describes phase oscillators on the unit circle whose interactions are encoded by a graph. Each edge acts like a spring that pulls the two adjacent oscillators toward each other whenever their phases differ. A central question is to determine which graphs are globally synchronizing, meaning that trajectories of the Kuramoto dynamics converge to the fully synchronized state from almost all initial conditions, except for a set of measure zero. This property is tightly linked to the benign nonconvexity of the model’s energy landscape. Existing guarantees for global synchronization rely on minimum-degree thresholds, which require the graph to be highly dense. In this work, we show that connected threshold graphs, whose density may vary from $2/n$ to $1$, are globally synchronizing. Our proof relies on a phasor-geometric analysis of the stationary points of the associated energy landscape.
@@ 1. Introduction
The study of synchronization traces back to the 17th century, when Christiaan Huygens first observed the spontaneous synchronization of pendulum clocks. Since then, synchronization phenomena have been widely studied across both natural and technological systems. For example, the flashing of fireflies [Buck1968, Strogatz1993], the rhythmic beating of heart cells [Winfree1967, Mikhailov2002], synchronization in power grids [dorfler2012synchronization], coordinated motion in robotic networks [OlfatiSaber2006], modern artificial intelligence models such as neural networks [KuramotoNeurons2024], and Transformer architecture [Geshkovski2025, Criscitiello2024].
A major obstacle in the study of synchronization was the absence of a model that simultaneously offered analytical tractability and sufficient dynamical richness. The Kuramoto model, introduced by Yoshiki Kuramoto in 1975 [Kuramoto1975], provided precisely such a balance, and has since become a central and widely accepted model for analyzing synchronization phenomena.
The Kuramoto model describes $n$ oscillators on the unit circle, coupled according to a graph $G=(V,E)$ with adjacency matrix $A$. Mathematically, it is a set of ordinary differential equations
$\frac{d\theta_{i}}{dt}=\omega_{i}+\sum_{j=1}^{n}A_{ij}\sin(\theta_{j}-\theta_{i}),\qquad i\in V,$ (1)
where $\theta_{i}:\mathbb{R}\to\mathbb{S}^{1}$ denotes the phase of oscillator $i$. Specifically, each oscillator is driven by its intrinsic frequency $\omega_{i}$ and by the nonlinear interactions determined by the phase differences with its neighbors.
Early studies of the Kuramoto model primarily considered systems of oscillators with heterogeneous intrinsic frequencies $\{\omega_{i}\}_{i=1}^{n}$, coupled through a complete graph with uniform weights $\boldsymbol{A}_{ij}=K/n$. Here, $K$ denotes the strength of the coupling between oscillators. A fundamental question in this setting is to determine the critical coupling strength $K_{c}$, such that for $K>K_{c}$, the oscillators converge to a common frequency, a phenomenon known as frequency synchronization [^1] That is, all derivatives $\tfrac{d\theta_{i}(t)}{dt}$ become identical as $t\to\infty$., and for $K<K_{c}$, they remain desynchronized. This problem is of central interest across many scientific domains, ranging from biological synchronization, rhythmic phenomena [michaels1987sinoatrial], technological applications such as deep brain stimulation [tass2003desynchronizing, nabi2011singleinput] and power systems [chiang1995direct, dorfler2012synchronization]. For comprehensive surveys of this line of work and its applications, we refer to [strogatz2000, acebron2005, dorfler2014].
Another line of research focuses on the homogeneous Kuramoto model where all intrinsic frequencies are identical, i.e., $\omega_{i}=\omega$ for all $i\in\{1,\cdots,n\}$. In this case, frequency synchronization is always achieved. This becomes immediate after applying a co-rotating frame $\theta_{i}(t)\leftarrow\theta_{i}(t)-\omega t$, under which the dynamics reduce to the gradient system
$\frac{d\theta_{i}}{dt}=\sum_{j=1}^{n}A_{ij}\sin(\theta_{j}-\theta_{i}),\qquad i\in[n].$ (2)
By LaSalle’s invariance principle, every trajectory converges to the invariant set of equilibria satisfying $\dot{\theta}_{i}=0$ for all $i$. Thus, in the co-rotating frame, the asymptotic frequency of every oscillator is zero. Transforming back to the original coordinates, this implies that the oscillators always achieve frequency synchronization. In contrast, when intrinsic frequencies are non-identical, the dynamics may exhibit significantly more complex long-term behavior, including chaotic attractors [Maistrenko2005].
In the homogeneous setting, the remaining challenge is to determine whether phase synchronization occurs. This clearly depends on the initial conditions. If all oscillators start within a half circle, they converge to a common phase. However, beyond the half-circle region, the global behavior remained poorly understood for a long time.
Over the past two decades, a central question that has emerged is the following: under what graph structures does the homogeneous Kuramoto system achieve phase synchronization for all initial conditions, except for a set of measure zero? This problem is not only mathematically appealing but also closely connected to several recent developments across different research areas, including the Burer–Monteiro approach for semidefinite programming [bandeira2016low, ling2019landscape, ling2023local] and the dynamics of Transformer architectures [Geshkovski2025, geshkovski2023emergence].
In this work, we follow this line of research, and beginning with a precise formulation of the problem.
Problem Setting.
@@
We say that a graph $G=(V,E)$ is globally synchronizing if, for almost all initial conditions $\boldsymbol{\theta}(0)\in\mathbb{R}^{n}$, the solution to (2) satisfies
$\lim_{t\to\infty}(\theta_{i}(t)-\theta_{j}(t))=0\qquad\text{for all }i,j\in V.$
@@
Which graphs are globally synchronizing?
To build some intuition for the model (2) and the definition (1.1), one may imagine placing $n$ oscillators on the unit circle and connecting them by edges specified by the adjacency matrix $A$. If there is an edge between nodes $i$ and $j$, then whenever their phases differ, the interaction tends to reduce this difference and pull them toward a common phase.
[…]
Figure 1. A spring analogy for a coupled oscillator network. The label on each curved arrow marks which node exerts the force.
From the above illustration of the complete graph $K_{3}$, one may ask whether the three oscillators will eventually converge to a common position when placed as shown in Figure 1. Each node is subject to two forces, for instance, node $1$ is pulled toward node $2$ on one side and toward node $3$ on the other. So it is not immediately clear in which direction it will move, let alone whether all three nodes will ultimately coincide. Even in this simple example, the difficulty of the problem becomes apparent. Nevertheless, one may form an intuitive if somewhat heuristic expectation: since in a complete graph each oscillator tends to move closer to every other oscillator, the abundance of interactions may indeed drive the system to synchronize from almost all initial configurations.
Historically, Verwoerd and Mason [verwoerd2007global] conjectured in 2007 that every complete graph $K_{n}$ ($n\geq 2$) is globally synchronizing. This conjecture was resolved by Richard Taylor in 2012 [taylor2012there], and global synchronization of complete graphs has since become a standard benchmark result in this line of research. For classical results on the Kuramoto model on complete graphs, we refer to [acebron2005, strogatz2000, PikovskyRosenblum2015, Bick2020].
In the same work [taylor2012there], Taylor also initiated what is now known as the extremal combinatorics version of the problem [bandeira2025randomstrasse101openproblems2024], which sparked substantial interest and subsequent developments. A more detailed discussion of these advances will be given in Prior Work.
One standard approach to this problem is to study the corresponding energy landscape. Note that the Kuramoto model (2) can be formulated as a gradient flow of a certain energy function. Precisely, the system of ordinary differential equations (2) defines a gradient dynamical system (or called gradient flow [^2] We call a dynamical system a gradient flow on $E(x)$ if it has the form $\frac{dx(t)}{dt}=-\nabla E(x)$. Its trajectories evolve along the steepest descent direction of $E$.) with energy function $E_{G}(\boldsymbol{\theta})$, satisfying
$\frac{d\boldsymbol{\theta}}{dt}=-\nabla E_{G}(\boldsymbol{\theta})$ (3)
where $E_{G}(\boldsymbol{\theta}):(\mathbb{S}^{1})^{|V|}\rightarrow\mathbb{R}$ with graph $G=(V,E)$ is defined as
$E_{G}(\boldsymbol{\theta})$= $\frac{1}{2}\sum_{1\leq i,j\leq n}\boldsymbol{A}_{ij}\big(1-\cos(\theta_{i}-\theta_{j})\big)$ (4)
$=$ $\frac{1}{2}\sum_{1\leq i,j\leq n}\boldsymbol{A}_{ij}\big(1-\langle v_{i},v_{j}\rangle\big).$
Here, $v_{i}$ is the unit vector $(\cos\theta_{i},\sin\theta_{i})\in\mathbb{R}^{2}$.
@@
Clearly, this function is nonconvex. Moreover, understanding the geometry of (4) provides direct insight into the long-term dynamics of (2). In particular, the synchronous states correspond exactly to the global minima of this energy function, and those minima that are local but not global (called spurious local minimua) correspond to non-synchronous stable equilibria. If such spurious local minima exist, the gradient dynamics may converge to them whenever the initialization lies within their basin of attraction, in which case the graph $G$ is not globally synchronizing. In fact, if one can show that the energy function (4) admits no spurious second-order stationary point (in such case we say the function is benign), then the graph $G$ is globally synchronizing. A rigorous proof of this statement can be established using the Łojasiewicz gradient inequality [Lageman2007] together with the center manifold theorem; see Lemma A.1 in [Geshkovski2025] for an excellent exposition.
Prior Work. Existing guarantees for global synchronization on deterministic graphs focus on the extremal combinatorics version of the problem. The aim is to identify minimum-degree thresholds above which a graph is guaranteed to be globally synchronizing.
Problem 1.
Let $\delta(G)$ denote the minimum degree of $G$. Find a constant $\mu_{c}\in[0,1]$ such that
$\delta(G)\ \geq\ \mu_{c}(n-1)\ \Longrightarrow\ G\ \text{is globally synchronizing},$
whereas for every $\varepsilon>0$ there exist graphs with $\delta(G)\leq(\mu_{c}-\varepsilon)(n-1)$ that are not globally synchronizing.
The intuition behind this problem is that when every node is connected to a sufficiently large fraction of the network, the graph becomes dense enough for the collective attractive forces to guarantee convergence to synchrony. A remarkable property of these degree-based criteria is that, although they are only sufficient, they apply any graph meeting the threshold. This means global synchronization is guaranteed regardless of the graph’s finer structural details.
Several advances have been made. The first result, due to Taylor [taylor2012there] in 2012, established a minimum-degree threshold of 0.9375(n1)0.9375(n-1). Subsequently, Ling [ling2019landscape] proved that global synchronization holds whenever the minimum degree exceeds 0.7929(n1)0.7929(n-1), and Lu [lu2021synchronization] further improved this bound to 0.7889(n1)0.7889(n-1). The conjecture asserts that the minimum-degree threshold for global synchronization is exactly $3/4$, and below this density, even very dense graphs may fail to synchronize. We refer to Conjecture 5 (density threshold for global synchrony) in [bandeira2025randomstrasse101openproblems2024] for a formal statement of this problem.
Global synchrony on random graphs constitutes another important line of inquiry. Ling, Xu, and Bandeira [ling2019landscape] proved that for Erdős–Rényi graphs, global synchronization occurs with high probability as $n\to\infty$ provided that the edge probability satisfies $p\gg\log(n)/n^{1/3}$. This bound was later improved by Kassabov, Strogatz, and Townsend to $p\gg\log^{2}(n)/n$ [kassabov2022global]. Both works raised the conjecture that synchronization should already occur as soon as the graph is connected, that is, for any $\epsilon>0$ and $p\geq(1+\epsilon)\log(n)/n$. This conjecture was recently resolved in [abdalla2024expandergraphsgloballysynchronizing]. Further improvement has been made in [jain2025random, mcrae2025benign]. For results on global synchrony of random geometric graphs, see [abdalla2024guarantees].
Interestingly, recent work has pointed out that the idealized transformer dynamics admits a similar optimization form as the Kuramoto energy function [Geshkovski2025]. This gives rise to the so-called nonlinear setting of the Kuramoto model, obtained by placing a nonlinear function on top of the inner product $E_{f}(\boldsymbol{\theta})=\-\sum_{1\leq i,j\leq n}A_{ij}f\big(\cos(\theta_{i}-\theta_{j})\big).$ The classical Kuramoto case corresponds to the linear function $f(t)=t$, while in the transformer dynamics one takes $f(t)=e^{\beta t}$ where $\beta>0.$ In this setting, the central question is to determine the values of $\beta$ for which global synchronization occurs, that is, the tokens collapse into a single cluster. Equivalently, this means that all token are making uniform predictions. A well-known conjecture predicted that global synchronization holds for all $\beta\geq 0$. The problem was originally formulated in [Geshkovski2025], with an initial result for $\beta\geq\Omega(1/n)$. It was subsequently improved in [Criscitiello2024] to cover the range $\beta\in[0,1]$, and was very recently settled by Polyanskiy, Rigollet, and Yao [polyanskiy2025synchronization]. For more work connecting the Kuramoto model and Transformer (or attention) dynamics, see [geshkovski2023emergence, karagodin2024clustering, chen2025quantitative, polyanskiy2025synchronization, rigollet2025mean, Geshkovski2024] and the references therein.
Our Contribution. Existing sufficient conditions for global synchronization of deterministic graphs require the graph to be sufficiently dense. In this work, we present a class of graphs that does not satisfy this minimum-degree requirement and whose densities range from 0 to $1$, yet globally synchronizes nevertheless. Through this result. Our main result is the following:
@@
In the following, we briefly introduce the definition of threshold graphs and the motivation for studying their global synchronization. We conclude this section with a discussion of the structural mechanism that underlies the global synchronization of this graph class.
(1) What Are Threshold Graphs? Threshold graphs can be constructed starting from a single vertex by iteratively adding each new vertex as either: an isolated vertex connecting to no existing vertex, or a dominating vertex connecting to all existing vertices. This generative process is uniquely encoded by a binary sequence of length n1n-1, where 0 denotes an isolated vertex and 1 a dominating one. Note that the binary sequence of any connected threshold graph must end with a 1; otherwise, the graph is not connected since there exists at least one isolated node. A full formal definition and additional characterizations are given in Section 2.
Figure 2. Four connected threshold graphs on 1919 vertices, from the sparsest ($S_{18}$) to the densest ($K_{19}$). Node labels indicate the construction order: vertex $1$ is the initial white vertex, and all subsequent labels follow the binary code of the construction. White and black circles denote isolated and dominating vertices (bit 0 and bit 11). Codes from left to right: 000000000000000001, 000100000000010001, 010001101101010001, $111111111$.
(2) Why Threshold Graphs? Many theorems begin with initial observations from which suitable conjectures are formulated, and our case is no exception. The investigation into global synchronization on threshold graphs began with several simple observations:
Observation 1. Star graphs and complete graphs are respectively the densest and sparsest connected threshold graphs, and both exhibit global synchronization.
$\bigl(\underbrace{\texttt{0}\;\texttt{0}\;\cdots\;\texttt{0}}_{\text{all 0}},\;\texttt{1}\bigr)$ $\bigl(\underbrace{\texttt{1}\;\texttt{1}\;\cdots\;\texttt{1}}_{\text{all 1}}\bigr)$
Star graph $S_{n}$ Complete graph $K_{n}$
Certainly, this is not sufficient to propose the conjecture that all threshold graphs globally synchronize. Indeed, although the two extreme cases synchronize, this does not by itself imply that the entire spectrum of threshold graphs does. However, two further observations strengthened our confidence in this conjecture.
Observation 2. One of the many equivalent definitions of threshold graphs requires that their neighborhood-inclusion preorder is total, i.e., the neighborhoods of any two nodes are nested: for any pair $u,v$, either $N(u)\subseteq N[v]$ or $N(v)\subseteq N(u)$. This total preorder of vicinities [threshold_graphs_book] suggests that influence may propagate step by step from the most dominant node to the rest without interference, hinting at the possibility of global synchronization.
Observation 3. While far from exhaustive, our numerical experiments did not reveal any non-synchronous stable equilibria on connected threshold graphs. Our experiments are twofold: (a) To examine global synchronization precisely in small systems, we employed computational algebraic geometry following [mehta2015algebraic]. More explicitly, we reformulated the equilibrium equations of the Kuramoto model as a system of polynomial equations and used the Bertini software package [BHSW13] to compute all solutions. We then evaluated the Jacobian at each solution to determine its stability. (b) For larger size systems (up to 200), using ode45 in Matlab, we initialized a large number of random initial conditions and observed that all trajectories converged to the fully synchronized state.
(3) Differences Between Our Approach and Existing Techniques. Many existing proofs of global synchronization rely, either explicitly or implicitly, on the so-called half-circle lemma. This lemma states that if all oscillators are confined to an open half-circle, then their dynamics will contract toward synchrony. Based on this observation, the bootstrap strategy attempts to show that all second-order stationary points must lie within a half-circle. If successful, this rules out the existence of non-synchronous state that the dynamics may converge to and thereby implies global synchronization.
@@
In this work, we bypass the half-circle approach by reducing the high-dimensional geometric question of whether the energy landscape is benign to a planar vector-geometric condition. This condition naturally leads us to study twins in the graph-theoretic sense, as well as a broader notion that we introduce, which we call geometric twins. The latter depends jointly on the graph structure and on the equilibrium configuration $\boldsymbol{\theta}$. These notions clarify how the relative positions of phases behave at second-order stationary point of the energy landscape $E(\boldsymbol{\theta})$. Building on this understanding, we employ an inductive strategy: synchronization is first established locally on small substructures, and then shown to propagate across the whole graph along the neighborhood-inclusion hierarchy of threshold graphs.
(4) Understanding the Mechanism Behind Global Synchrony in Threshold Graphs. Previous global synchronization results focus on density-based conditions. For instance, any graph whose minimum degree exceeds 0.7889(n1)0.7889(n-1) is globally synchronizing [lu2020synchronization]. In other words, these results require the graph to be sufficiently dense. Threshold graphs, however, lie outside this regime. Depending on the positions of dominating vertices in the generating sequence, their density can vary from $O(2/n)$ (stars) to $1$ (cliques), and many intermediate densities occur. Nevertheless, we prove in this work that they also exhibit global synchronization. What deeper insight can we draw from this result?
Our analysis shows that the global synchronization of threshold graphs is not a consequence of density. Instead, it arises from a different mechanism, namely the presence of local symmetries. Some of these symmetries come from the graph structure itself: as we show in Lemma 5.2, closed twins must synchronize at any second-order stationary point. There are also symmetries that do not originate purely from combinatorial structure, that is, certain configurations exhibit a form of geometric symmetry. As we will show in Section 5.3, in the configurations we call synchronous pendants, pairs of vertices that are not structurally symmetric nonetheless become geometrically equivalent at second-order stationary points and hence synchronize. Such symmetric configurations arise inductively along the construction process of threshold graphs. Their synchronization propagates backward from later-added vertices to earlier ones, which shows that every second-order stationary point is a synchronous state. Consequently, every threshold graph is globally synchronizing.
@@
Notation. Throughout this work, we consider only connected simple graphs, i.e., graphs without self-loops or weighted edges. Following the convention of dynamical systems, we refer to the variable $\theta=(\theta_{1},\cdots,\theta_{n})$ of the Kuramoto model (2) as a state, and call the solution $\boldsymbol{\theta}(t)$ with respect to the initial condition $\boldsymbol{\theta}(0)$ in the Kuramoto model (2) a trajectory starting from $\boldsymbol{\theta}(0)$. Moreover, $\mathbb{S}^{1}:=\{x\in\mathbb{R}^{2}:\|x\|=1\}$ denotes the unit circle. And $\boldsymbol{x}\Uparrow\boldsymbol{y}$ means that $\boldsymbol{x}$ and $\boldsymbol{y}$ are nonzero and positively collinear: $\boldsymbol{x}=\lambda\boldsymbol{y}$ for some $\lambda>0$. We write $u\sim v$ to indicate that vertices $u$ and $v$ are adjacent in a graph, and $u\nsim v$ otherwise. For a vertex $i$ and a subset $Q$ of vertices, we denote
$N_{Q}(i)=\{\,q\in Q:q\sim i\,\},$
the set of neighbors of $i$ that lie in $Q$.